Given n positive numbber in real number in R the product of their harmonic mean and their arithemetic mean is 1

Grade:12th pass

1 Answers

Rohit pandey
6 years ago
let, a & b be two real positive numbers. Then , A.M of a & b = a + b. H.M = a*b/(a+b).Thus, A.M*H.M=a*b.Now, according to the question , a*b= 1.i.e; a = 1/b.Therefore, two positive real numbers ben & 1/ n.